OpenStudy (anonymous):

factor 2x^2 – 5x – 12?

OpenStudy (anonymous):

(2x +3 )(x -4 )

OpenStudy (anonymous):

do u need any explanation?

OpenStudy (anonymous):

1) Find roots of 2x^2 – 5x – 12 = 0 2) Answer is: 2(x-1st root)(x-2nd root)

OpenStudy (anonymous):

Yeah an explanation would be good.

OpenStudy (anonymous):

well the 2x and x in the brackets are easy - this gives the 2x^2 now we need to get 2 numbers which will give you the middle value -5x - the two numbers must also give -12 when multiplied together and 1 of these 2 numbers must also multiply 2 to give the -5. its now trial and error :- 3 x -4 gives -12 and we can get -5 if we multiply the -4 by the 2 (from the 2x^2). -4 x 2 + 3 = 15 its a good idea to put them into columns as below 2x^2 -12 -5x ----------------------------- 2 x 1 4 x -3 2x 4 - 3 = 5 - no 2 x 1 -4x3 2 x-4 + 3 = -5 yes

OpenStudy (anonymous):

thats a bit long winded i know - but practice theses a lot and it will come to you quite easily

OpenStudy (anonymous):

error in the above i meant to write -4 x 2 - 3 = -5